# Accurate estimation of feature importance faithfulness for tree models

### v1

@@ Appendix A Models training
The grid search was performed by using class $GridSearchCV$ from library XGBoost [6] with numbers of folds in cross-validation set to $5$. Additionally to parameters described in 4.2, the grid search included also parameters with the following possible values:
@@
Appendix B Additional experiments results
B.1 Results from precision experiments
@@
B.2 Results from PGI comparison experiment
Figure 11: Single tree model for the Wine Quality dataset.
Figure 12: Bigger model for the Wine Quality dataset.
Figure 13: Single tree model for Californian Housing dataset
Figure 14: Bigger model for Californian Housing dataset

### v3

@@ Appendix A Models training
The grid search was performed by using class $GridSearchCV$ from library XGBoost (Chen and Guestrin 2016) to minimize squared error, with numbers of folds in cross-validation set to $5$. Additionally to parameters described in Models, the grid search included also parameters with the following possible values:
  • $max\_depth$ with possible values in $\{1,2,3,4\}$
@@
The chosen parameters for models were:
  • Single tree model for Wine Quality dataset – $eta=0.9$, $subsample=0.9$and $max\_depth=4$
  • Bigger model for Wine Quality dataset – $eta=0.2$, $subsample=0.9$, $max\_depth=4$ and $40$ trees
  • Single tree model for Californian Housing dataset – $eta=0.9$, $subsample=0.9$ and $max\_depth=3$
  • Bigger model for Californian Housing dataset – $eta=0.2$, $subsample=0.9$, $max\_depth=4$ $40$ trees
  • Single tree model for Parkinson Telemonitoring dataset – $eta=0.9$ and $subsample=0.9$, $max\_depth=3$
  • Bigger model for Parkinson Telemonitoring dataset – $eta=0.3$, $subsample=0.8$ and $max\_depth=4$
Appendix B Quasi Monte Carlo
In our implementation of the Quasi-Monte Carlo method, we utilized the quasi-random generator provided by the SciPy library (Virtanen et al. 2020). Specifically, we employed the Halton sequence (Halton 1960) for generating our quasi-random points. This choice was motivated by the Halton sequence’s flexibility, as it does not necessitate a sample size that is a power of two $(2^{n})$, being a better fit for experimental designs. To transform the uniform distribution produced by the Halton sequence into the required normal distribution, we applied the inverse transform method. It is worth noting that while this transformation is effective, it does result in some loss of the low-discrepancy properties inherent to the original Halton sequence.
Appendix C Additional experiments results
Results from precision experiments
@@ Appendix C Additional experiments results
Figure 6: Single tree model for Parkinson Telemonitoring dataset
Figure 7: Bigger model for Parkinson Telemonitoring dataset
Results from feature removing experiment
Table 4: Rankings comparison wrt. feature randomising metric
$k$ SHAP $\sigma$ = $0.1$ $\sigma$ = $0.3$ $\sigma$ = $1.0$
Bigger Red Wine Quality model
1 0.944 0.922 0.923 0.919
2 0.893 0.902 0.881 0.891
3 0.858 0.89 0.87 0.868
Bigger California Housing model
1 1.32e+05 1.412e+05 1.379e+05 1.264e+05
2 1.3e+05 1.354e+05 1.321e+05 1.263e+05
3 1.282e+05 1.318e+05 1.287e+05 1.27e+05
Bigger Parkinson Telemonitoring model
1 8.87 8.96 8.99 8.945
2 8.748 8.882 8.882 8.906
3 8.585 8.777 8.809 8.784
Table 5: Rankings comparison wrt. remove-and-retrain metric
$k$ SHAP $\sigma$ = $0.1$ $\sigma$ = $0.3$ $\sigma$ = $1.0$
Bigger PGI Red Wine Quality model
1 0.67 0.621 0.637 0.667
2 0.72 0.638 0.66 0.672
3 0.722 0.652 0.665 0.711
Bigger PGI California Housing model
1 8.092e+04 7.187e+04 7.206e+04 7.485e+04
2 8.913e+04 7.693e+04 8.062e+04 8.707e+04
3 9.136e+04 8.267e+04 8.965e+04 9.665e+04
Bigger PGI Parkinson Telemonitoring model
1 7.216 6.745 6.863 7.02
2 7.388 6.951 7.138 7.292
3 7.55 7.021 7.282 7.45
Results from PGI comparison experiment
Table 6: Greedy $\mathrm{PG}^{2}$ vs SHAP rankings wrt. $\mathrm{PGI}^{2}$.
$\sigma^{\prime}$ SHAP $\sigma$ = $0.1$ $\sigma$ = $0.3$ $\sigma$ = $1.0$
Single Red Wine Quality model
0.1 0.021 0.023 0.023 0.023
0.3 0.077 0.095 0.096 0.096
1.0 0.151 0.174 0.185 0.185
Single California Housing model
0.1 3.267e+08 3.987e+08 3.987e+08 3.987e+08
0.3 1.421e+09 1.582e+09 1.597e+09 1.597e+09
1.0 5.921e+09 6.276e+09 6.451e+09 6.459e+09
Single Parkinson Telemonitoring model
0.1 1.084 1.084 1.084 1.084
0.3 3.438 3.573 3.823 3.823
1.0 6.143 7.322 8.444 8.449